# Optimizing Personalized Federated Learning through Adaptive Layer-Wise Learning

### v1

@@ Abstract
Real-life deployment of federated Learning (FL) often faces non-IID data, which leads to poor accuracy and slow convergence. Personalized FL (pFL) tackles these issues by tailoring local models to individual data sources and using weighted aggregation methods for client-specific learning. However, existing pFL methods often fail to provide each local model with global knowledge on demand while maintaining low computational overhead. Additionally, local models tend to over-personalize their data during the training process, potentially dropping previously acquired global information. We propose FLAYER, a novel layer-wise learning method for pFL that optimizes local model personalization performance. FLAYER considers the different roles and learning abilities of neural network layers of individual local models. It incorporates global information for each local model as needed to initialize the local model cost-effectively. It then dynamically adjusts learning rates for each layer during local training, optimizing the personalized learning process for each local model while preserving global knowledge. Additionally, to enhance global representation in pFL, FLAYER selectively uploads parameters for global aggregation in a layer-wise manner. We evaluate FLAYER on four representative datasets in computer vision and natural language processing domains. Compared to six state-of-the-art pFL methods, FLAYER improves the inference accuracy, on average, by 5.42% (up to 14.29%).
@@ Introduction
Methods Dir(0.1) Dir(0.01) Pathological 20 clients 50 clients 100 clients Acc. Imps.
FedAvg 37.08$\pm$0.43 28.74$\pm$0.38 33.30$\pm$0.17 37.08$\pm$0.43 34.56$\pm$0.25 33.08$\pm$0.41 60.68$\pm$0.42 24.77
APPLE 57.29$\pm$0.30 74.52$\pm$0.19 67.69$\pm$0.10 57.29$\pm$0.30 58.09$\pm$0.24 48.46$\pm$0.32 - -
Ditto 48.46$\pm$0.89 72.94$\pm$0.22 63.98$\pm$0.09 48.46$\pm$0.89 46.08$\pm$0.19 43.42$\pm$0.37 58.49$\pm$0.21 10.03
FedAMP 48.75$\pm$0.49 73.12$\pm$0.17 63.91$\pm$0.32 48.75$\pm$0.49 46.49$\pm$0.44 43.74$\pm$0.20 60.72$\pm$0.27 11.97
FedPer 54.26$\pm$0.43 73.52$\pm$0.15 67.12$\pm$0.35 54.26$\pm$0.43 51.24$\pm$0.39 47.67$\pm$0.36 63.13$\pm$0.23 8.87
FedRep 53.94$\pm$0.40 75.08$\pm$0.18 68.01$\pm$0.30 53.94$\pm$0.40 50.10$\pm$0.30 45.80$\pm$0.27 61.33$\pm$0.17 7.39
FedALA 58.65$\pm$0.26 75.24$\pm$0.11 70.55$\pm$0.07 58.65$\pm$0.26 59.46$\pm$0.23 58.80$\pm$0.41 63.55$\pm$0.58 4.90
FLAYER #60.68$\pm$0.42 #77.39$\pm$0.24 #71.35$\pm$0.19 #60.68$\pm$0.42 #61.70$\pm$0.30 #59.96$\pm$0.39 - -
@@
We also evaluate the impact of statistical heterogeneity on FLAYER and other SOTA pFL methods using 20 clients. Specifically, we set two degrees of heterogeneity on CIFAR-100. The first scenario is $\beta=0.01$, where the smaller the value of $\beta$, the greater the heterogeneity of the setting. The second is an extreme case of heterogeneity, known as pathological heterogeneous setting (McMahan et al. 2017; Shamsian et al. 2021), where we sample 10 classes from the total 100 classes in CIFAR-100 for each client. We use $\beta=0.1$ as the baseline performance. Table 5 reports the performance impact under these varying degrees of heterogeneity. FLAYER consistently outperforms all other SOTA pFL methods across all heterogeneous settings, delivering the highest accuracy.

### v2

@@ Abstract
Real-life deployment of federated Learning (FL) often faces non-IID data, which leads to poor accuracy and slow convergence. Personalized FL (pFL) tackles these issues by tailoring local models to individual data sources and using weighted aggregation methods for client-specific learning. However, existing pFL methods often fail to provide each local model with global knowledge on demand while maintaining low computational overhead. Additionally, local models tend to over-personalize their data during the training process, potentially dropping previously acquired global information. We propose FLAYER, a novel layer-wise learning method for pFL that optimizes local model personalization performance. FLAYER considers the different roles and learning abilities of neural network layers of individual local models. It incorporates global information for each local model as needed to initialize the local model cost-effectively. It then dynamically adjusts learning rates for each layer during local training, optimizing the personalized learning process for each local model while preserving global knowledge. Additionally, to enhance global representation in pFL, FLAYER selectively uploads parameters for global aggregation in a layer-wise manner. We evaluate FLAYER on four representative datasets in computer vision and natural language processing domains. Compared to six state-of-the-art pFL methods, FLAYER improves the inference accuracy, on average, by 7.21% (up to 14.29%).
@@ Introduction
Methods Dir(0.1) Dir(0.01) 20 clients 50 clients 100 clients Acc. Imps.
FedAvg 37.08$\pm$0.43 43.74$\pm$0.38 37.08$\pm$0.43 34.56$\pm$0.25 33.08$\pm$0.41 60.68$\pm$0.42 24.77
APPLE 57.29$\pm$0.30 74.52$\pm$0.19 57.29$\pm$0.30 58.09$\pm$0.24 48.46$\pm$0.32 - -
Ditto 48.46$\pm$0.89 72.94$\pm$0.22 48.46$\pm$0.89 46.08$\pm$0.19 43.42$\pm$0.37 58.49$\pm$0.21 10.03
FedAMP 48.75$\pm$0.49 73.12$\pm$0.17 48.75$\pm$0.49 46.49$\pm$0.44 43.74$\pm$0.20 60.72$\pm$0.27 11.97
GPFL 51.06$\pm$0.42 74.59$\pm$0.21 51.06$\pm$0.42 48.30$\pm$0.29 44.61$\pm$0.32 - -
FedPer 54.26$\pm$0.43 73.52$\pm$0.15 54.26$\pm$0.43 51.24$\pm$0.39 47.67$\pm$0.36 63.13$\pm$0.23 8.87
FedRep 53.94$\pm$0.40 75.08$\pm$0.18 53.94$\pm$0.40 50.10$\pm$0.30 45.80$\pm$0.27 61.33$\pm$0.17 7.39
FedCP 46.72$\pm$0.38 69.42$\pm$0.32 46.72$\pm$0.38 42.86$\pm$0.33 40.19$\pm$0.24 - -
FedALA 58.65$\pm$0.26 75.24$\pm$0.11 58.65$\pm$0.26 59.46$\pm$0.23 58.80$\pm$0.41 63.55$\pm$0.58 4.90
FLAYER #60.68$\pm$0.42 #77.39$\pm$0.24 #60.68$\pm$0.42 #61.70$\pm$0.30 #59.96$\pm$0.39 - -
@@
We also evaluate the impact of statistical heterogeneity on FLAYER and other SOTA pFL methods using 20 clients. Specifically, we set two degrees of heterogeneity on CIFAR-100. The first scenario is $\beta=0.01$, where the smaller the value of $\beta$, the greater the heterogeneity of the setting. We use $\beta=0.1$ as the baseline performance. Table 5 reports the performance impact under these varying degrees of heterogeneity. FLAYER consistently outperforms all other SOTA pFL methods across all heterogeneous settings, delivering the highest accuracy.